# Counterfactual Fairness by Combining Factual and Counterfactual Predictions

### v1

@@ 4 Related Works
Fairness Notions Fair Machine Learning has accumulated a vast literature that proposes various notions to measure fairness issues of machine learning models. Representative fairness notions can be categorized into three classes. Group fairness such as demographic parity (Pedreshi et al., 2008) and equalized odds (Hardt et al., 2016) requires certain group-level statistical independence between model predictions and individual’s demographic information. Despite its conceptual simplicity, group fairness is known for ruling out perfect model performance (Hardt et al., 2016) and may allow for bias against certain individuals (Corbett-Davies et al., 2023). Individual fairness (Dwork et al., 2012), on the other hand, asks a model to treat similar individuals similarly. However, it is often a highly task-specific and open-ended question how to measure the similarity between different individuals. Recently, counterfactual fairness (CF, Kusner et al. (2017)) further takes the causal relationship of data attributes into consideration when measuring fairness. In words, counterfactual fairness proposes that a model should treat any individual same to their counterfactual if the individual had been from another demographic group. As an individual-level notion agnostic to the choice of similarity measure (Kusner et al., 2017), CF has recently gained traction (Wu et al., 2019; Nilforoshan et al., 2022; Makhlouf et al., 2022; Rosenblatt and Witter, 2023b). Motivated by these recent advances, in this work we focus on the counterfactual fairness.
@@
Methods for Fairness Given an unfair dataset, attempts to achieve fairness fall into three categories. Pre-processing cleans the data before running machine learning models on it by resampling the samples or removing undesired attributes (Kamiran and Calders, 2012). In-processing intervenes the model-training process by incorporating fairness constraints (Zafar et al., 2017; Donini et al., 2018; Lohaus et al., 2020) or penalties (Mohler et al., 2018; Scutari et al., 2021; Liu et al., 2023). Post-processing adjusts the raw model outputs to close the bias gap by, e.g., assigning each demographic group a unique decision threshold (Jang et al., 2022). Post-processing has been favored as a more efficient and practical solution that does not require retraining the original model (Petersen et al., 2021; Xian et al., 2023). To achieve CF, Kusner et al. (2017) applied pre-processing and discarded all descendants of the sensitive feature. Garg et al. (2019); Stefano et al. (2020); Kim et al. (2021) in-processed the model training by penalizing CF violations but their solutions lack formal CF guarantees and often contain unsatisfactory bias after the intervention (Zuo et al., 2023). Recently, Zuo et al. (2023) proposed another in-processing based solution that is capable of achieving perfect CF and better performance. However, their model still suffered from a performance drops, and it is unknown whether their method is optimal.
@@
Inherent Trade-off between Fairness and Predictiveness Machine learning models are known suffered from performance drops after fairness interventions (Hardt et al., 2016; Menon and Williamson, 2018; Chen et al., 2018), which is known as the fairness-utility trade-offs. Recently, inherent trade-offs towards non-causal based fairness such as demographic parity (DP) has been established separately for regression Chzhen et al. (2020) and classification tasks Xian et al. (2023). The excess risks are characterized by certain distribution distance (i.e., Wasserstein-2 barycenter for regression, and total-variation or Wasserstein-1 barycenter for noiseless or noisy classification) between the conditional distribution of $Y$ given $A$. A similar trade-off between CF and predictiveness has also been empirically observed (Zuo et al., 2023). Nonetheless, their inherent trade-off remains an open open question. In this work we take the first step towards this goal and provide a quantitative analysis in both complete and incomplete causal knowledge settings as presented in Section 3. We hope our work sheds lights on future works towards more effective CF.
@@ 5 Experiments
In this section, we validate our theorems and effectiveness of our algorithms through experiments on synthetic datasets and semi-synthetic datasets. On synthetic datasets, we focus on validating our theorems in the settings where our assumptions are met. On semi-synthetic datasets, we focus on validating the effectiveness of our methods in more practical scenarios where limited causal knowledge is available and the invertibility assumption is relaxed.
@@
We consider two metrics in this paper: Error and Total Effect (TE). The former evaluates if each method can achieve its goal regardless of fair or not. This is important because we can achieve perfect Counterfactual Fairness by always outputing fixed prediction given whatever input, but that is not useful at all. The latter is a common metric to evaluate Counterfactual Fairness (Kim et al., 2021; Zuo et al., 2023). Given a test set $\mathcal{D}_{\textnormal{test}}$, Error is defined as $\textnormal{Error}=\frac{1}{|\mathcal{D}_{\textnormal{test}}|}\sum_{x^{(i)}\in\mathcal{D}_{\textnormal{test}}}\ell(\widehat{y}(x^{(i)}),y^{(i)})$ where $y^{(i)}$ is the ground truth target, $\widehat{y}(x^{(i)})$ is the prediction of $x^{(i)}$, and $\ell$ depends on the task. TE is defined as $\textnormal{TE}=\frac{1}{|\mathcal{D}_{\textnormal{test}}|}\sum_{x^{(i)}\in\mathcal{D}_{\textnormal{test}}}|\widehat{y}(x^{(i)})-\widehat{y}(x_{a^{\prime}}^{(i)})|$ where $x_{a^{\prime}}$ is the ground truth counterfactual corresponding to $x^{(i)}$. Since we only consider binary sensitive attribute, we further define $\textnormal{TE}_{0}=\frac{1}{|\{i:a^{(i)}=0\}|}\sum_{i:a^{(i)}=0}|\widehat{y}(x^{(i)})-\widehat{y}(x_{a^{\prime}}^{(i)})|$ and $\textnormal{TE}_{1}=\frac{1}{|\{i:a^{(i)}=1\}|}\sum_{i:a^{(i)}=1}|\widehat{y}(x^{(i)})-\widehat{y}(x_{a^{\prime}}^{(i)})|$ to evaluate Counterfactual Fairness for different group respectively.
@@ Methods
In general, we consider the following methods: (1) Empirical Risk Minimization (ERM): Directly train a classifier on all features without any fainress consideration. Specifically $\widehat{y}=\phi(x,a),$ where $\phi$ represents the predictor. (2) Counterfactual Fairness with $U$ (CFU) (Kusner et al., 2017): To achieve Counterfactual Fairness, CFU proposes to use $U$ for prediction. Specifically, $\widehat{y}=\phi(u)$. (3) Counterfatual Fairness with fair representation (CFR) (Zuo et al., 2023): CFR proposes to use $U$ and a symmetric version of $x,x_{a^{\prime}}$. Specifically, $\widehat{y}=\phi(\frac{x+x_{a^{\prime}}}{2},u)$. (4) PCF [^3] Essentially PCF with ERM (PCF-ERM). For brevity, we just call it PCF.: As introduced in Algorithm 1, PCF proposes to mix the output of factual and counterfactual prediction. Specifically, $\hat{y}=p(a)\phi(x,a)+(1-p(a))\phi(x_{a}^{\prime},a^{\prime})$. (5) PCF with analytic solution (PCF-Ana): In our synthetic experiment, instead of training via ERM, we can directly acquire bayes optimal $\phi$ in closed-form. Detailed can be found in Section B.2. (6) PCF with CRM (PCF-CRM): As discussed in Section 3.3, it could be hard to get bayes optimal predictor, especially when there is counterfactual estimation error. Here due to the scale of our experiment, we choose to augment the dataset with estimated counterfactuals rather than finetuning. Specifically, $\phi$ is trained via ERM on the dataset $\mathcal{D}_{\textnormal{train}}=\{x^{(i)},y^{(i)},a^{(i)}\}_{i=1}^{N}\cup\{\hat{x}^{(i)}_{1-a},y^{(i)},1-a^{(i)}\}_{i=1}^{N}$.
@@ 5.1 Synthetic Dataset
The classification tasks take the same form except $Y\sim\textnormal{Bernoulli}(\sigma(w_{X}X+w^{\prime}_{U}U+w_{Y}\epsilon_{Y})$ and $Y\sim\textnormal{Bernoulli}(\sigma(w_{X}X^{3}+w^{\prime}_{U}U+w_{Y}\epsilon_{Y}))$ for Linear-Cls and Cubic-Cls respectively. More details could be found in Section B.1. Results are averaged over 5 different runs where the structural model is kept the same but data is resampled. All results shown in the main paper use KNN based predictor. Results with other predictors can be found in Appendix C.
@@
Figure 3: Results on synthetic datasets under counterfactual estimation error. Different color represents different $\alpha$ indicating the standard deviation of the error ($\epsilon\sim\mathcal{N}(0,\alpha)$) while shape represents different algorithms. Results with different $\beta$ can be found in Appendix C.
@@
Figure 4: Results on synthetic datasets comparing PCF and PCF-Analytic. Different color represents different $\alpha$ indicating the standard deviation of the error ($\epsilon\sim\mathcal{N}(0,\alpha)$) while shape represents different algorithms. Results with different $\beta$ can be found in Appendix C.
@@ 5.2 Semi-synthetic Dataset
We consider Law School Success dataset (Wightman, 1998) in this section. The main goal of this experiment is to validate the effectiveness of our methods in more practical scenarios where limited causal knowledge is available and the invertibility assumption is relaxed.
To compute TE, we need access to ground truth counterfactuals. Hence we train a generative model on real dataset to generate semi-synthetic dataset following the method in Zuo et al. (2023). Note that the counterfactuals are hidden from the algorithm and just used for evaluation. More details could be found in Section B.1. All experiments are repeated 5 times on the same semi-synthetic dataset.
@@ Results
In Figure 5, we observe that PCF-CRM achieves better CF and lower Error in comparison to CFU and CFR. This validates our improvement Section 3.3.2 indeed leads to more practical algorithm. Results on comparing PCF and PCF-CRM can be found in Appendix C, which further justifies this. While ERM could achieve lower error, it has worst fairness. This is inevitably determined by the inherent trade-off discussed in Theorem 3.4. Furthermore, inspired by the trade-off, we test the result of mixing all predictors with ERM. The curve shows that PCF-CRM remains optimal given fixed CF and best CF given fixed error. This demonstrates again that PCF-CRM is the best among all methods.
@@ Conclusion
In this work, we conducted a formal study on the trade-off between Counterfactual Fairness (CF) and predictive performance. We proved that combining factual and counterfactual predictions with a potentially unfair optimal predictor can achieve optimal CF. We also proved the excess risk between predictors with and without CF constraints, offering insights into the minimal predictiveness degradation needed for perfect CF. To address incomplete causal knowledge, we examined the impact of imperfect counterfactual estimations on CF and predictive performance. We proposed a plugin method using pre-trained models for optimal fair prediction and a practical approach to mitigate estimation inaccuracies.
Despite the theoretical merits of our method, there are two limitations that might raise concerns about its practical applicability: (1) access to ground truth counterfactuals and (2) access to Bayes optimal predictors. Here we discuss these limitations in more depth to bolster understanding of how our methods can be used in practice and how it can benefit from the broader community. Hopefully, this discussion will also inspire future research directions.
@@
While how to better estimate counterfactuals is out of the scope of this work, it is indeed an unavoidable challenge faced by the community of Counterfactual Fairness. It not only limits the deployment of CF algorithms, but also leads to difficulty in validating proposed CF methods. There are some works in the field of causality that aims at estimating counterfactuals. For instance, Nasr-Esfahany et al. (2023) proves counterfactual identifiablity under certain causal graphs. However, in more general scenarios, we sometimes lack such knowledge, and identifying the causal graph itself can be challenging. These tasks have been well studied in the field of causal discovery (Chickering, 2002; Colombo et al., 2014) and causal representation learning (Schölkopf et al., 2021). Solutions to this problem typically rely on strong assumptions, such as the linearity of Structural Causal Models (SCMs) or additive noise (Shimizu et al., 2006; Hoyer et al., 2008; Peters et al., 2014). More recently, Zhou et al. (2024) proposes a method of estimating counterfactuals without need to identify the causal model or graph, which we believe have the potential to be a good plugin counnterfactual estimator in our algorithm.

### v2

@@ 4 Related Works
Fairness Notions Fair Machine Learning has accumulated a vast literature that proposes various notions to measure fairness issues of machine learning models. Representative fairness notions can be categorized into three classes. Group fairness, such as demographic parity (Pedreshi et al., 2008) and equalized odds (Hardt et al., 2016), requires certain group-level statistical independence between model predictions and individuals’ demographic information. Despite its conceptual simplicity, group fairness is known for ruling out perfect model performance (Hardt et al., 2016) and may allow for bias against certain individuals (Corbett-Davies et al., 2023). Individual fairness (Dwork et al., 2012), on the other hand, asks a model to treat similar individuals similarly. However, determining the similarity between different individuals is often highly task-specific and open-ended. Recently, counterfactual fairness (CF, Kusner et al. (2017)) further takes the causal relationship of data attributes into consideration when measuring fairness. In words, counterfactual fairness proposes that a model should treat any individual the same as their counterfactual if the individual had been from another demographic group. As an individual-level notion agnostic to the choice of similarity measure (Kusner et al., 2017), CF has recently gained traction (Wu et al., 2019; Nilforoshan et al., 2022; Makhlouf et al., 2022; Rosenblatt and Witter, 2023). Motivated by these recent advances, in this work, we focus on the counterfactual fairness.
@@
Methods for Fairness Given an unfair dataset, attempts to achieve fairness fall into three categories. Pre-processing cleans the data before running machine learning models on it, typically by resampling samples or removing undesired attributes (Kamiran and Calders, 2012). In-processing intervenes the model-training process by incorporating fairness constraints (Zafar et al., 2017; Donini et al., 2018; Lohaus et al., 2020) or penalties (Mohler et al., 2018; Scutari et al., 2021; Liu et al., 2023). Post-processing adjusts the raw model outputs to close the bias gap by, e.g., assigning each demographic group a unique decision threshold (Jang et al., 2022). Post-processing has been favored as an efficient and practical solution because it does not require retraining the original model (Petersen et al., 2021; Xian et al., 2023). To achieve CF, Kusner et al. (2017) applied pre-processing and discarded all descendants of the sensitive feature. Chen et al. (2024) pre-processed the data via orthogonalization and marginal distribution mapping. Garg et al. (2019); Stefano et al. (2020); Kim et al. (2021) in-processed the model training by penalizing CF violations but their solutions lack formal CF guarantees and often contain unsatisfactory bias after the intervention (Zuo et al., 2023). Recently, Zuo et al. (2023) proposed another in-processing based solution that is capable of achieving perfect CF and better performance. However, their model still suffered performance drops, and it is unknown whether their method is optimal. Wang et al. (2023) leveraged predictor that satisfies equal counterfactual opportunity criterion to construct a counterfactually fair predictor. While they provide results on optimality, their findings assume an ideal setting where non-sensitive features are independent of sensitive features.
@@
Inherent Trade-off between Fairness and Predictiveness Machine learning models are known to suffer from performance drops after fairness interventions (Hardt et al., 2016; Menon and Williamson, 2018; Chen et al., 2018), which is known as the fairness-utility trade-offs. Recently, inherent trade-offs towards non-causal based fairness such as demographic parity (DP) has been established separately for regression (Chzhen et al., 2020) and classification tasks (Xian et al., 2023). The excess risks are characterized by certain distribution distance (i.e., Wasserstein-2 barycenter for regression, and total-variation or Wasserstein-1 barycenter for noiseless or noisy classification) between the conditional distribution of $Y$ given $A$. A similar trade-off between CF and predictiveness has also been empirically observed (Zuo et al., 2023). Nonetheless, their inherent trade-off remains an open question. In this work we take the first step towards this goal and provide a quantitative analysis in both complete and incomplete causal knowledge settings as presented in Section 3. We hope our work sheds light on future works towards more effective CF.
@@ 5 Experiments
In this section, we validate our theorems and the effectiveness of our algorithms through experiments on synthetic and semi-synthetic datasets. On synthetic datasets, we focus on validating our theorems in settings where our assumptions hold. On semi-synthetic datasets, we aim to assess the effectiveness of our methods in more practical scenarios, where limited causal knowledge is available and the invertibility assumption is relaxed.
@@
We consider two metrics in this paper: Error and Total Effect (TE). The former evaluates whether each method can achieve its goal, irrespective of fairness. This is important because we can achieve perfect Counterfactual Fairness by always outputting fixed prediction given whatever input, but that is not useful at all. The latter is a common metric to evaluate Counterfactual Fairness (Kim et al., 2021; Zuo et al., 2023). Given a test set $\mathcal{D}_{\textnormal{test}}$, Error is defined as $\textnormal{Error}=\frac{1}{|\mathcal{D}_{\textnormal{test}}|}\sum_{x^{(i)}\in\mathcal{D}_{\textnormal{test}}}\ell(\widehat{y}(x^{(i)}),y^{(i)})$ where $y^{(i)}$ is the ground truth target, $\widehat{y}(x^{(i)})$ is the prediction of $x^{(i)}$, and $\ell$ depends on the task. TE is defined as $\textnormal{TE}=\frac{1}{|\mathcal{D}_{\textnormal{test}}|}\sum_{x^{(i)}\in\mathcal{D}_{\textnormal{test}}}|\widehat{y}(x^{(i)})-\widehat{y}(x_{1-a}^{(i)})|$ where $x_{1-a}$ is the ground truth counterfactual corresponding to $x^{(i)}$. Since we only consider binary sensitive attribute, we further define $\textnormal{TE}_{0}=\frac{1}{|\{i:a^{(i)}=0\}|}\sum_{i:a^{(i)}=0}|\widehat{y}(x^{(i)})-\widehat{y}(x_{1-a}^{(i)})|$ and $\textnormal{TE}_{1}=\frac{1}{|\{i:a^{(i)}=1\}|}\sum_{i:a^{(i)}=1}|\widehat{y}(x^{(i)})-\widehat{y}(x_{1-a}^{(i)})|$ to evaluate Counterfactual Fairness for different group respectively.
@@ Methods
In general, we consider the following methods: (1) Empirical Risk Minimization (ERM): Directly train a classifier on all features without any fairness consideration. Specifically $\widehat{y}=\phi(x,a),$ where $\phi$ represents the predictor. (2) Counterfactual Fairness with $U$ (CFU) (Kusner et al., 2017): To achieve Counterfactual Fairness, CFU proposes to use $U$ for prediction. Specifically, $\widehat{y}=\phi(u)$. (3) Counterfactual Fairness with fair representation (CFR) (Zuo et al., 2023): CFR proposes to use $U$ and a symmetric version of $x,x_{1-a}$. Specifically, $\widehat{y}=\phi(\frac{x+x_{1-a}}{2},u)$. (4) Equal Counterfactual Opportunity (ECOCF)(Wang et al., 2023): An ECO predictor is adjusted to become counterfactually fair. Specifically, $\hat{y}=p(a)[p(a)\phi(x,a)+(1-p(a))\phi(x,1-a)]+(1-p(a))[(1-p(a))\phi(x_{1-a},1-a)+p(a)\phi(x_{1-a},a)]$ (5) PCF [^4] Essentially PCF with ERM (PCF-ERM). For brevity, we just call it PCF.: As introduced in Algorithm 1, PCF proposes to mix the output of factual and counterfactual prediction. Specifically, $\hat{y}=p(a)\phi(x,a)+(1-p(a))\phi(x_{1-a},1-a)$. (6) PCF with analytic solution (PCF-Ana): In our synthetic experiment, instead of training via ERM, we can directly acquire bayes optimal $\phi$ in closed-form. Detailed can be found in Section C.2. (7) PCF with CRM (PCF-CRM): As discussed in Section 3.3, it could be hard to get the optimal predictor when there is counterfactual estimation error. Here due to the scale of our experiment, we augment the dataset with estimated counterfactuals rather than finetuning. Specifically, $\phi$ is trained via ERM on the dataset $\mathcal{D}_{\textnormal{train}}=\{x^{(i)},y^{(i)},a^{(i)}\}_{i=1}^{N}\cup\{\hat{x}^{(i)}_{1-a},y^{(i)},1-a^{(i)}\}_{i=1}^{N}$.
@@ 5.1 Synthetic Dataset
The classification tasks take the same form except $Y\sim\textnormal{Bernoulli}(\sigma(w_{X}X+w^{\prime}_{U}U+w_{Y}\epsilon_{Y})$ and $Y\sim\textnormal{Bernoulli}(\sigma(w_{X}X^{3}+w^{\prime}_{U}U+w_{Y}\epsilon_{Y}))$ for Linear-Cls and Cubic-Cls respectively. More details could be found in Section C.1. Results are averaged over 5 different runs where the structural model is kept the same but data is resampled. All results shown in the main paper use KNN based predictor. Results with other predictors can be found in Appendix D.
@@
Figure 3: Results on synthetic datasets under counterfactual estimation error. Different color represents different $\alpha$ indicating the standard deviation of the error ($\epsilon\sim\mathcal{N}(0,\alpha)$) while shape represents different algorithms. Results with different $\beta$ can be found in Appendix D.
@@
Figure 4: Results on synthetic datasets comparing PCF and PCF-Analytic. Different color represents different $\alpha$ indicating the standard deviation of the error ($\epsilon\sim\mathcal{N}(0,\alpha)$) while shape represents different algorithms. Results with different $\beta$ can be found in Appendix D.
@@
Figure 5: Results on Sim-Law with estimated counterfatuals. The predictor is a MLP regressor. We also test the convex combination of each algorithm and ERM. For example, PCF-CRM with $\lambda$ means $\hat{y}=\lambda\hat{y}_{\textnormal{PCF-CRM}}+(1-\lambda)\hat{y}_{\textnormal{ERM}}$. This suggests that PCF-CRM can achieve lower Error given the same TE and lower TE given the same Error.
@@ 5.2 Semi-synthetic Dataset
In this section, we consider Law School Success dataset (Wightman, 1998) where the sensitive attribute is gender and the target is first-year grade.. The main goal of this experiment is to validate the effectiveness of our methods in more practical scenarios where limited causal knowledge is available and the invertibility assumption is relaxed.
To compute TE, we need access to ground truth counterfactuals. Hence we train a generative model on real dataset to generate semi-synthetic dataset following the method in Zuo et al. (2023). We want to emphasize that counterfactuals are hidden from downstream models and used for the evaluation of TE only. This way, we get access to the ground truth $u^{*}$ and can generate ground truth counterfactuals without any error. In our investigation, exogenous noise, factual data and counterfactual data are all actually the simulated version of original datasets. However they do follow a fixed data generating mechanism that is close to the real data. More details could be found in Section C.1. All experiments are repeated 5 times on the same semi-synthetic dataset.
@@ Results
In Figure 5, we observe that PCF-CRM achieves better CF and lower Error in comparison to CFU and CFR. This validates our improvement Section 3.3.2 indeed leads to more practical algorithm. Results on comparing PCF and PCF-CRM can be found in Appendix D, which further justifies this. While ERM could achieve lower error, it has worst fairness. This is inevitably determined by the inherent trade-off discussed in Theorem 3.4. Furthermore, inspired by the trade-off, we test the result of mixing all predictors with ERM. The curve shows that PCF-CRM remains optimal given fixed CF and best CF given fixed error. This demonstrates again that PCF-CRM is the best among all methods.
@@ Conclusion
In this work, we conducted a formal investigation of the trade-off between Counterfactual Fairness (CF) and predictive performance. We proved that combining factual and counterfactual predictions with a potentially unfair, optimal predictor achieves optimal CF. Additionally, we derived the excess risk between predictors with and without CF constraints, quantifying the minimum performance degradation necessary to ensure perfect CF. To address incomplete causal knowledge, we analyzed the effects of imperfect counterfactual estimations on CF and predictive performance. We proposed a plugin approach that leverages pre-trained models for optimal fair prediction and developed a practical method to mitigate estimation errors.
Despite the theoretical contributions of our method, two limitations may impact practical applicability: (1) access to ground-truth counterfactuals and (2) access to Bayes optimal predictors. Below, we delve into these limitations, clarifying how our methods can be practically applied and how they can benefit from contributions from the broader community. We hope this discussion will also inspire future research directions.
@@
While how to better estimate counterfactuals is out of the scope of this work, it is indeed an unavoidable challenge faced by the community of Counterfactual Fairness. It not only limits the deployment of CF algorithms, but also leads to difficulty in validating proposed CF methods. While counterfactual data can be obtained in specific scenarios, such as through randomized controlled trials, it is challenging to acquire in most applications. There are some works in the field of causality that aims at estimating counterfactuals. For instance, Nasr-Esfahany et al. (2023) proves counterfactual identifiability under certain causal graphs. However, in more general scenarios, such causal knowledge may be lacking and identifying the causal graph itself can be challenging. These tasks have been well studied in the field of causal discovery (Chickering, 2002; Colombo et al., 2014) and causal representation learning (Schölkopf et al., 2021). Solutions to this problem typically rely on strong assumptions, such as the linearity of Structural Causal Models (SCMs) or additive noise (Shimizu et al., 2006; Hoyer et al., 2008; Peters et al., 2014). More recently, Zhou et al. (2024) propose a method of estimating counterfactuals without the need to identify the causal model or graph. We believe this approach to direct counterfactual estimation could have the potential to be a good plugin counnterfactual estimator in our algorithm. Additionally, generative models could also be used to generate samples as if they had come from a different sensitive attribute (Choi et al., 2018; Zhou et al., 2022b, 2023; Rombach et al., 2022). These methods often offer the advantage of higher sample quality, especially in modalities such as images or natural language. However, they must be applied with considerable care, as they generally lack integration with the causal model and may introduce significant estimation errors.